The sum of a number, and the difference between 16 and a number, plus 5 is 21. What is the number
Answer by checkley71(8403)

X+(X-16)+5=21
X+X-16+5=21
2X=21+16-5
2X=32
X=32/2
X=16 SOLUTION
PROOF
16+(16-16)+5=21
16=5=21
21=21